# A Perspective for Adapting Generalist AI to Specialized Medical AI Applications and Their Challenges

### v1

@@ 1 Introduction
Artificial intelligence (AI) is increasingly being integrated into various medical tasks, including clinical risk prediction [1], medical image understanding [2], and synthetic patient records generation [3]. Typically, these models are designed for specific tasks and will struggle with unfamiliar tasks or out-of-distribution data [4]. Large language models (LLMs) are foundation AI models characterized by their extensive training data and enormous model scale. Unlike traditional AI models, LLMs demonstrate an emergent capability in language understanding and the ability to tackle new tasks through in-context learning [5]. For example, we can teach LLMs to conduct a new task by providing the text explanation of the task (or “prompts”), the input and output protocols, and several examples. This adaptability has sparked interest in employing generalist LLMs to medical AI applications such as chatbots for outpatient reception [6]. Contrary to the common belief that generalist LLMs will excel in many fields [7], we advocate that domain-specific AI adaptations for medicine are more effective and safe. This paper will overview how various adaptation strategies can be developed for medical AI applications and the associated trade-offs.
@@
Generalist LLMs such as ChatGPT can support broad tasks but may underperform in specialized domains [8]. One notable drawback is the occurrence of “hallucinations,” which are fabricated facts that look plausible yet incorrect [9]. High-stakes medical applications such as patient-facing diagnosis tools are especially vulnerable to such inaccurate information [10]. In response, adaptation methods in LLM for medicine have thrived, focusing on enhancing LLMs’ domain-specific capabilities (Fig. 1b). They include finetuning LLMs on medical data [11], adding relevant medical information to the prompts for LLMs via retrieval-augmented generation (RAG) [12], and equipping LLMs with external tools to building AI agents achieving autonomous planning and task execution [13]. With the increasing practice in developing LLM-based AI applications, it is becoming evident that cutting-edge performance is increasingly driven by the mixture of these adaptations [14] and systematic engineering of multiple AI components [15].
We present a systematic overview of adaptation methods and organize them in a framework for developing LLM-based medical AI applications. Fig. 1 outlines various adaptation methods and high-level strategies, including:
  • Model development: This includes techniques like pretraining, where medical knowledge is injected into generic models via continual pretraining on medical datasets [16], and finetuning, which ensures that the model’s outputs are consistent with domain knowledge and human preferences [17].
  • Model optimization: Strategies such as prompt optimization enhance the relevance and accuracy of model responses [14], while retrieval-augmentation generation (RAG) utilizes external data to create richer, more informed prompts [12]. More advanced operations include building autonomous AI agents [18] and chaining multiple AI modules to enable complex workflows [19].
  • System engineering: We can optimize the performance of medical AI by taking a system engineering approach. This includes subtasking by assigning clearly defined, narrow-scope tasks to each system component to improve reliability [20], involving human-in-the-loop to boost oversight and refinement [21], result verification for maintaining accuracy and robustness [22], and interconnected AI agents focusing different aspects of the task for improving medical conversational reasoning [23].
Next, we describe this framework for developing medical AI applications in detail, focusing on diverse medical AI use cases. We will end by describing the associated challenges and opportunities for further development of LLM-based medical AI applications.
@@
Figure 1: Workflow for AI model development, optimization, and system engineering in medical applications. (a) Generalist AI models, such as proprietary systems (e.g., GPT-4) and open-source models (e.g., LLaMA), serve as foundational technologies for developing specialized medical AI models. (b) Adapting generalist AI to medical tasks involves several techniques, including model fine-tuning, prompt optimization, and the development of AI agents or AI chains. These methods use diverse medical datasets, such as medical images, electronic health records (EHRs), clinical notes, publications, and omics data, to enhance AI model training and performance. (c) Effective system engineering for medical AI entails integrating AI modules into comprehensive chains to support tasks like cohort extraction, eligibility assessment, and result verification. This process emphasizes human interaction and AI, resulting in tailored AI modules for specific applications. (d) Generalist AI applications in medicine span various domains, including conversational diagnosis, radiology report generation, clinical note summarization, automated medical coding, drug design, patient-trial matching, and systematic literature reviews. All require advanced system integration for optimal performance.
@@ 2.1 Model development: building medical-specific LLMs
Pretraining
Generic LLMs such as ChatGPT are trained on vast, diverse datasets to develop a broad understanding of language [5]. A continual pretraining of generic LLMs on medical data, such as medical publications, clinical notes, and electronic health records (EHRs), can enhance LLMs’ alignment in medical languages. For example, MEDITRON is based on the open-source LLaMA model [24], further trained on medical data including PubMed articles and medical guidelines. It performs comparably to larger generalist models such as GPT-4 [25] in medical question answering. Another model, Panacea, shows the benefit of LLMs in clinical trial search, design, and patient recruitment via a continual pretraining of generic LLMs on clinical trial documents, including trial protocols and trial-related publications [16].
Instruction finetuning
Instead of pretaining with medical data without specific task supervision, an LLM can also learn from paired custom queries and their expected responses to perform multiple target tasks. This training process is known as instruction finetuning [26]. For example, when PaLM, a 540-billion parameter LLM, underwent instruction tuning, it resulted in MedPaLM, which achieved 67.6% accuracy on US Medical Licensing Exam-style questions [11].
@@ 2.2 Model optimization: strategies for improving LLM performance
Prompt indicates the user inputs to LLMs, which can include the task description, the expected input and output formats, and some input/output examples. For example, this input “Your task is to summarize the input clinical note. Please adhere to these summarization standards: […list of criteria]. Refer to these examples: […list of examples]. Keep in mind: […list of important hints],” can guide LLMs towards generating the summary of patient notes. This practice leverages the principle of in-context learning, where LLMs adapt to new tasks based on the input prompts without additional training [5]. Research shows that the structure and content of the prompt significantly affect the model’s performance. For instance, chain-of-thought prompting [28] encourages LLMs to engage in multiple steps of reasoning and self-reflection, thereby enhancing the output quality. Another strategy involves ensembling [29, 30], where outputs derived from multiple prompts are synthesized to produce a final, more robust response. In the medical domain, MedPrompt [14] has demonstrated its ability to outperform domain-tuned LLMs by combining multiple prompting techniques. Additionally, TrialGPT [31] effectively adapts LLMs to match patient notes to the eligibility criteria of a clinical trial through prompting.
@@
Handcrafted prompts are highly dependent on domain knowledge and trial-and-error, but they can still perform suboptimally and cause reliability issues in applications. Techniques such as automatic prompt generation [32] and optimization [33] were hence proposed. These approaches can transform the adaptation of LLMs for medical tasks into a more structured machine-learning task. For example, requesting an LLM to summarize clinical notes could start with a simple prompt like “Summarize the input clinical note.” Using a collection of example notes and expert-written summaries, automatic summarization evaluation metrics can serve as the target for iterative prompt refinement. The prompt can ultimately evolve into a more advanced prompt consisting of professional task description, representative examples, and clear output format requirement description.
Retrieval-augmented generation
@@
Retrieval-augmented generation (RAG) is an extension of prompting. It dynamically incorporates information retrieved from external databases into the model’s inputs, supplementing LLM’s internal knowledge [34]. In medicine, it is extensively employed to improve the factual accuracy of LLM responses to medical questions by fetching the evidence from medical literature, clinical guidelines, or medical ontology [12, 35]. In practice, the efficacy of RAG is critically dependent on the performance of the retrieval systems, which are responsible for identifying and extracting the most relevant content to supplement the LLM’s output [36, 37]. For the clinical notes summary task, the RAG strategy should first retrieve the most relevant clinical information from the patient’s records and then use it to construct the prompt dynamically.
AI agent
@@
An LLM-based AI agent embodies a vision where AI systems actively solve problems through autonomous planning, knowledge acquisition using diverse tools, and iterative self-reflection [38]. This concept is supported by various agent frameworks that facilitate these capabilities, such as AutoGPT and AutoGen [39, 40]. Beyond external data sources, these agents can interact with various external tools, enhancing their operational scope. This includes interfaces for programming [41], autonomous laboratory equipment [18], and integration with other software programs [42, 43], broadening the potential applications of AI agents. In medicine, AI agents are making significant strides. For instance, conversational agents are being developed for medical diagnosis, providing real-time patient interaction and support [44]. Additionally, agents are increasingly used in biomedical research, assisting in complex data analysis and hypothesis generation [45].
AI chain
@@
It is often challenging for LLMs to handle complex tasks requiring multiple steps of observation and reasoning in a single conversation [28]. To address this, instead of having LLMs split tasks and conduct them with several AI agents, experts can break down tasks into multiple steps. As such, multiple LLM calls form an AI chain, enhancing both transparency and controllability of the final output [46]. An illustrative example is WikiChat [47], which addresses hallucinations by sequentially employing LLMs to query Wikipedia, summarize, filter, and extract the retrieved facts, perform additional fact-checking, and finally draft responses with a round of self-reflection for refinement. As a result, WikiChat surpassed GPT-4 by 55% in factual accuracy for question-answer tasks. In addition to optimizing performance, FrugalGPT [48] optimizes query routing to reduce inference costs by using a cascade of LLMs with varying capabilities. This system utilizes weaker LLMs for initial processing and escalates to stronger LLMs only as necessary.
@@
2.3 System engineering: developing compound medical AI systems
In addition to breaking down tasks into multiple steps, building medical AI systems requires thoughtful design of modules to execute these substeps effectively. Equally important is the orchestration of these modules into a cohesive pipeline, a process we refer to as subtasking. We split the substaking development into two levels: modules and pipelines.
  • Modules are crafted using LLM operation tools (e.g., prompt optimization), typically featuring a specialized prompt and designed for interaction with other system components.
  • Pipelines, on the other hand, integrate these modules in either a cascaded or a more complex arrangement to optimize performance for specific medical tasks.
Module development
An AI module for medical applications is characterized by its objectives (e.g., progress note generation) and constraints (i.e., response time and token usage). It is task-specific and is designed with a protocol for its input-output format, ensuring seamless integration into a pipeline. There are several critical factors to be considered.
First, we should describe the primary objective function of the AI module. This function often relies on a metric that evaluates the quality of the outputs. For instance, when evaluating the patient summaries from EHRs, metrics such as completeness, correctness, and conciseness can be considered [51]. Additional secondary objectives may include metrics related to the rate of exceptions, computational cost, and response time. Moreover, a module may incorporate specific constraints, including limitations on time or token usage or requirements concerning the output’s length, format, and content. For example, TrialMind [52], an LLM-driven system for medical literature mining, consists of multiple specialized modules for different subtasks, including literature search, screening, and information extraction. The search module aims to improve the recall of relevant studies by refining search queries while working within constraints related to query length and complexity to manage the total number of retrieved studies. The screening module focuses on enhancing the precision of study eligibility assessment. It provides a list of eligibility scores for each input criterion in a strict format, which is then used to calculate an overall study-level eligibility score.
Second, we should identify and optimize various parameters of the AI module. A key parameter is the selection of the underlying LLM. Choices can be made between using proprietary LLMs like GPT-4 or fine-tuning open-source models like Llama [24]. Assessments of available LLMs should consider trade-offs in quality, speed, cost, and privacy. For example, developing and deploying a local model is recommended when handling highly sensitive patient records or clinical trial documents [16]. Another crucial parameter is the curated prompt. Various optimizations can be applied, such as incorporating demonstration examples of input sentences and output labels to enable few-shot clinical information extraction [53].
Third, external tools can complement the AI modules in medical applications. A common example is the retriever, which accesses domain-specific knowledge or datasets to support retrieval-augmented generation (RAG). For instance, GeneGPT [54] leverages biomedical databases from the National Center for Biotechnology Information [55] to enhance the factual accuracy of biomedical question answering, achieving an accuracy of 0.83 compared to ChatGPT’s 0.12. Additionally, integrating external programs to generate and execute code can improve LLMs’ performance in medical tasks requiring numerical reasoning, such as analyzing real-world patient data [56].
Pipeline development
A medical AI pipeline consists of multiple modules tailored to meet the quality and format requirements of the target medical task. Here are key elements for pipeline development. First, a medical task follows a structured, multi-step process that must adhere to established medical protocols or practices. Application developers must map out the necessary steps and orchestrate suitable modules for each phase to assemble the overall pipeline. For example, a medical systematic review usually follows the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guideline [57]. Referring to PRISMA, an AI pipeline can be crafted with modules for study search, study eligibility assessment, risk of bias assessment, study characteristic extraction, and results extraction [52].
Second, integrating multiple AI modules presents challenges in establishing cohesive communication protocols, as the output of one module becomes the input for another. This process requires software engineering practices, and using a programming language for AI modules could help standardize and streamline AI application development [58]. Langchain [59] and LLamaIndex [60], two widely-used frameworks, address this by compiling advanced prompting techniques into modular building blocks, simplifying LLM application development.
Result verification
The growing complexity of medical AI applications presents challenges in identifying and addressing failure modes. To address this, it is essential to incorporate verification steps into intermediate outputs, allowing for detecting failure sources and, when possible, automatic correction. DSPy [20], a programming language that abstracts AI applications into a computational pipeline, offers a solution by introducing a compiler that automatically generates optimized strategies for invoking LLMs. Users can embed assertions into the pipeline to enforce format constraints and enable automatic self-correction, thereby enhancing the reliability of AI applications. For example, DSPy has been successfully used to detect and correct errors in clinical texts [19].
Human in the loop
Current LLM outputs may not be completely stable or trustworthy. When used for important medical tasks, close human supervision is necessary [61]. This requirement encourages the design of user interfaces for effective human-in-the-loop integration [62]. This interface may encompass actions such as chatting, selecting, and clicking. The interface’s nature, like whether it supports text, images, code, or a combination of modalities, not only shapes the user experience but also influences the selection and design of the pipeline modules. For example, a system for medical image-based diagnosis requires an interface where users can upload pathology images. The outputs should provide the diagnosis and reference images from the database, illustrating how the AI arrived at its prediction [63]. A copilot for structured medical data analysis needs to display multiple coding steps and their intermediate outputs, allowing experts to instantly verify and modify the synthesized programs [41].

### v3

@@ 1 Introduction
Artificial intelligence (AI) is increasingly being integrated into various medical tasks, including clinical risk prediction [1], medical image understanding [2], and synthetic patient records generation [3]. Typically, these models are designed for specific tasks and will struggle with unfamiliar tasks or out-of-distribution data [4]. Large language models (LLMs) are foundation models characterized by their extensive training data and enormous model scale. Unlike traditional AI models, LLMs demonstrate an emergent capability in language understanding and the ability to tackle new tasks through in-context learning [5]. For example, we can teach LLMs to conduct a new task by providing the text explanation of the task (or “prompts”), the input and output protocols, and several examples. This adaptability has sparked interest in employing generalist LLMs to medical AI applications such as chatbots for outpatient reception [6]. Contrary to the common belief that generalist LLMs will excel in many fields [7], we advocate that domain-specific AI adaptations for medicine are more effective and safe. This paper will overview how various adaptation strategies can be developed for medical AI applications and the associated trade-offs.
@@
Generalist LLMs such as ChatGPT can support broad tasks but may underperform in specialized domains [8]. One notable drawback is the occurrence of “hallucinations,” which are fabricated facts that look plausible yet incorrect [9]. High-stakes medical applications such as patient-facing diagnosis tools are especially vulnerable to such inaccurate information [10]. In response, adaptation methods in LLM for medicine have thrived, focusing on enhancing LLMs’ domain-specific capabilities (Fig. 1b). They include finetuning LLMs on medical data [11], adding relevant medical information to the prompts for LLMs via retrieval-augmented generation (RAG) [12, 13], and equipping LLMs with external tools to building AI agents achieving autonomous planning and task execution [14, 15]. With the increasing practice in developing LLM-based AI applications, it is becoming evident that cutting-edge performance is increasingly driven by the mixture of these adaptations [16] and systematic engineering of multiple AI components [17].
In this Perspective, we present an overview of adaptation techniques for developing LLM-based medical AI and the workflow to solidify the development (Fig. 1). The adaptations can be concluded as:
  • Model development focuses on designing model architectures and employing learning algorithms to adapt model parameters for medical tasks. Techniques include injecting medical knowledge into general-purpose models through continual pretraining on medical datasets [18] and fine-tuning, which aligns the model’s outputs with domain-specific knowledge and human preferences [19].
  • Model optimization enhances model performance by optimizing its associated components, such as optimizing the input prompts [16] and implementing retrievers accessing external data to enable retrieval-augmented generation [12].
  • System engineering enhances medical AI performance by breaking down tasks into well-defined, narrow-scope components. LLMs can serve as the computational core for each specialized component, which can be linked together to support complex workflows [20], or interact autonomously with other components to form agent-based systems [21].
Next, we provide actionable guidelines on when and how to adopt these adaptation methods based on the specific task parameters, e.g., time and cost constraints. Additionally, we present concrete use cases that demonstrate the practical value of this framework. Finally, we discuss the associated challenges and opportunities for advancing LLM-based medical AI applications.
@@
Figure 1: Workflow for adapting generalist LLMs for medical AI through adaptation techniques. (a) Generalist AI models, such as proprietary systems (e.g., Open AI’s GPT-4) and open-source models (e.g., LLaMA), serve as foundational technologies for developing specialized medical AI models. (b) Adapting generalist AI to medical tasks involves several techniques, including model fine-tuning, prompt optimization, and the development of AI agents or AI chains. These methods use diverse medical datasets, such as medical images, electronic health records (EHRs), clinical notes, publications, and omics data, to enhance AI model training and performance. (c) Effective system engineering for medical AI entails integrating AI modules into comprehensive chains to support tasks like cohort extraction, eligibility assessment, and result verification. This process emphasizes human interaction and AI, resulting in tailored AI modules for specific applications. (d) Generalist AI applications in medicine span various domains, including conversational diagnosis, radiology report generation, clinical note summarization, automated medical coding, drug design, patient-trial matching, and systematic literature reviews. All require advanced system integration for optimal performance.
@@ 2.1 Model development: building medical-specific LLMs
Generic LLMs such as ChatGPT benefit from vast parameters and are trained on vast, diverse datasets to develop a broad understanding of language [5]. A continual pretraining of generic LLMs on medical data, such as medical publications, clinical notes, and electronic health records (EHRs), can enhance LLMs’ alignment in medical languages. For example, MEDITRON is based on the open-source LLaMA model [22], further trained on medical data including PubMed articles and medical guidelines. It performs comparably to larger generalist models such as GPT-4 [23] in medical question answering. Another model, PANACEA, shows the benefit of LLMs in clinical trial search, design, and patient recruitment via a continual pretraining of generic LLMs on clinical trial documents, including trial protocols and trial-related publications [18]. In addition, scaling model size using a mixture-of-experts strategy, where multiple LLMs (i.e., experts) are combined with a router network to select the appropriate expert during inference, has been found to yield superior performance at significantly lower computational cost compared to using a single, larger LLM [24].
Instead of pretaining with medical data without specific task supervision, an LLM can also learn from paired custom queries and their expected responses to perform multiple target tasks. A widely used technique is instruction tuning [25]. For example, when PaLM, a 540-billion parameter LLM, underwent instruction tuning, it resulted in MedPaLM, which achieved 67.6% accuracy on US Medical Licensing Exam-style questions [11]. When the computational resources are limited, we can also update a subset of the model’s parameters or attach an additional small but learnable component to the model (e.g., prefix-tuning [26] and LoRA [27]).
@@ 2.2 Model optimization: strategies for improving LLM performance
Prompt indicates the inputs to LLMs, which can include the task description, the expected input and output formats, and some input/output examples. For example, this input “Your task is to summarize the input clinical note. Please adhere to these summarization standards: […list of criteria]. Refer to these examples: […list of examples]. Keep in mind: […list of important hints],” can guide LLMs towards generating the summary of patient notes. This practice leverages the principle of in-context learning, where LLMs adapt to new tasks based on the input prompts without additional training [5]. Research shows that the structure and content of the prompt significantly affect the model’s performance. For instance, chain-of-thought prompting [29] encourages LLMs to engage in multiple steps of reasoning and self-reflection, thereby enhancing the output quality. Another strategy involves ensembling [30, 31], where outputs derived from multiple prompts are synthesized to produce a final, more robust response. In the medical domain, MedPrompt [16] has demonstrated its ability to outperform domain-tuned LLMs by combining multiple prompting techniques. Additionally, TrialGPT [32] effectively adapts LLMs to match patient notes to the eligibility criteria of a clinical trial through prompting.
@@
Handcrafted prompts are highly dependent on domain knowledge and trial-and-error, but they can still perform suboptimally and cause reliability issues in applications. Techniques such as automatic prompt generation [33] and optimization [34] were proposed. These approaches can transform the adaptation of LLMs for medical tasks into a more structured machine-learning task. For example, requesting an LLM to summarize clinical notes could start with a simple prompt like “Summarize the input clinical note.” Using a collection of example notes and expert-written summaries, automatic summarization evaluation metrics can serve as the target for iterative prompt refinement. The prompt can ultimately evolve into a more advanced prompt consisting of a professional task description, representative examples, and a clear output format requirement description.
@@
Retrieval-augmented generation (RAG) is an extension of prompting. It dynamically incorporates information retrieved from external databases into the model’s inputs, supplementing LLM’s internal knowledge [12]. In medicine, it is extensively employed to improve the factual accuracy of LLM responses to medical questions by fetching the evidence from medical literature, clinical guidelines, or medical ontology [13, 35]. In practice, the efficacy of RAG relies on two aspects: (1) the quality of the external database, which should be high-quality and up-to-date [36, 37] and (2) the performance of the retrieval systems, which are responsible for identifying and extracting the most relevant content to supplement the LLM’s output [38, 39].
Figure 2: This playbook outlines the process of adapting large language models (LLMs) for medical AI using a systems engineering approach. (a) Selecting the overall architecture of the system should be based on the properties and requirements of the task at hand. (b) When building agent systems, four main modules need to be developed. LLMs act in different roles when equipped with different modules and interact with human experts to dynamically conduct the target task. (c) When building AI chain systems, we can first define the pipeline that decomposes the task into small steps following expert workflow or professional guidelines and then develop the module responsible for each step. (d) Adaptation techniques can be applied to enhance LLM’s performance for the AI agent or for the AI module. Adaptation methods need to be selected according to data availability and task requirements.
2.3 System engineering: developing compound medical AI systems
Fig. 2 illustrates the top-down system engineering process for adapting LLMs to medical AI applications. The process involves the following steps: (1) selecting the architecture to be developed (Fig. 2a); (2) designing the agent-based system (Fig. 2b) or the AI chain system (Fig. 2c); and (3) developing and fine-tuning LLMs to integrate with these systems (Fig. 2d).
Architecture selection
The first step is to choose the overall system architecture: AI Chain [40] and AI Agent [41], depending on the nature of the tasks and requirements, as illustrated in Fig. 2a. AI Chain is suitable for fixed, routinely structured workflows, making it ideal for repetitive tasks, professional practices, and scenarios where adherence to guidelines or compliance is critical. Example applications include systematic literature reviews [42] and patient-to-trial matching [32], where processes are well-defined and consistent. On the other hand, AI agents are better suited for flexible workflows and exploratory tasks that are less sensitive to time, cost, and uncertainty. This paradigm is ideal for dynamic and creative tasks such as data science code generation [43] and brainstorming research hypotheses [44]. The choice of architecture ensures that the system is aligned with the specific operational demands and objectives of the intended medical AI application.
Agent system development
@@ 2.3 System engineering: developing compound medical AI systems
Developing AI agent systems boils down to designing and integrating key modules to facilitate the interactions between LLMs and human experts, as demonstrated in Fig. 2b. AI agents embody a vision where AI systems actively solve problems through autonomous planning, knowledge acquisition using diverse tools, and iterative self-reflection [41]. This concept is supported by various agent frameworks that facilitate these capabilities, such as AutoGPT and AutoGen [45, 46]. Beyond external data sources, these agents can interact with various external tools, enhancing their operational scope. This includes interfaces for programming [47], autonomous laboratory equipment [21], and integration with other software programs [48, 49], broadening the potential applications of AI agents.
In general, an AI agent is composed of a base generalist AI like LLM and one or multiple modules so as to have different roles in conducting the task collaboratively with human experts. These modules include Reasoning, which supports both direct reasoning and reasoning with feedback from humans, external tools, or other agents, to enhance decision-making capabilities; Perception, which enables the agent to process and understand various data modalities such as text, images, and code; Memory, comprising internal memory for storing temporary context (e.g., in prompts) and external memory for long-term information storage (e.g., in vector databases); and Interaction, which supports communication with external tools, coordination with multiple agents, and collaboration between humans and AI.
In medicine, AI agents are making remarkable progress in transforming healthcare delivery. For example, Polaris is an agent system designed to facilitate patient-facing, real-time healthcare conversations [50]. It integrates a large language model (LLM) with automatic speech recognition for speech-to-text transcription and text-to-speech technology to convert LLM-generated text into audio, enabling seamless real-time interactions with patients. By leveraging agent-specific prompts and coordinating with specialized sub-AI models, such as those focused on medication management, laboratory and vital sign analysis, and clinical nutrition, Polaris functions as a multidisciplinary virtual specialist. Additionally, agents are increasingly used in biomedical research, assisting in complex data analysis and hypothesis generation [51].
AI chain development
@@
For tasks that are usually fixed, repetitive, or strictly governed by professional guidelines, the AI chain is recommended to be developed (Fig. 2c). Rather than allowing LLMs to divide tasks and conduct them with several AI agents, experts can break down tasks into a chain of multiple steps. This approach allows multiple LLM calls to create an AI chain, which enhances both the transparency and controllability of the final output [40]. An illustrative example is WikiChat [52], which addresses hallucinations by sequentially employing LLMs to query Wikipedia, summarize, filter, and extract the retrieved facts, perform additional fact-checking, and finally draft responses with a round of self-reflection for refinement. As a result, WikiChat surpassed GPT-4 by 55% in factual accuracy for question-answer tasks. In addition to optimizing performance, FrugalGPT [53] optimizes query routing to reduce inference costs by using a cascade of LLMs with varying capabilities. This system utilizes weaker LLMs for initial processing and escalates to stronger LLMs only as necessary.
AI chain development, as shown in Fig. 2c, consists of three primary stages: pipeline development, module definition, and module development. The process begins with designing the pipeline based on expert workflows or professional guidelines, breaking it down into subtasks such as search, reasoning, and data extraction. Next, specialized modules are developed for each subtask. Each module is defined by its objectives (e.g., progress note generation) and constraints (e.g., response time and token usage) and is optimized using adjustable parameters such as the underlying models, prompts, and external tools or data sources. During module development, the focus is on the underlying model adaptations, considering the specifics of module inputs and outputs, including their format (e.g., text, image, code), length, and frequency (e.g., real-time or offline batch processing). For instance, the input length is an important factor for a module for summarizing multiple long documents, which requires an underlying model with a long context length optimized for improving fidelity to its inputs [54]. For even longer inputs, advanced adaptations such as map-reduce, i.e., segmenting inputs, summarizing each segment, and consolidating those summaries into a global summary [55]. This process also requires orchestrating data flows, integrating inputs from upstream modules, external components (e.g., knowledge bases, vector databases, or machine learning models), and human requests. The outputs need to undergo automatic or human verification to ensure accuracy and reliability.
@@
Adaptation techniques
The next step is the adaptation of LLMs, which includes two stages: model selection and model development and optimization, illustrated in Fig. 2d. In the model selection stage, the choice between proprietary and open-source models is guided by specific application requirements. Proprietary models, such as ChatGPT, are accessed via cloud services, enabling lightweight development and deployment without the need for local GPU resources but often at a higher cost, making them suitable for well-funded organizations. In contrast, open-source models such as finetuned LLaMa models offer greater flexibility for customization and are more cost-effective for high-volume tasks. They are particularly advantageous for local deployment scenarios due to privacy or security compliance requirements.
The second stage emphasizes optimizing the selected model based on data availability and task requirements. Prompt optimization is the recommended initial approach for cases with limited or no labeled data, particularly when working with proprietary models [58]. However, if sufficient domain-specific data is available, continual pretraining can be applied to inject domain knowledge into generalist LLMs, enhancing their adaptability to specialized tasks. Nonetheless, sometimes the gain may be marginal if the domain-specific data are publicly available because training data for generic LLMs may already include the domain-specific data [59]. When prompt optimization fails to meet performance expectations, we recommend collecting more task demonstrations (i.e., input-output pairs) and finetuning the generalist models [60]. Finetuning offers several advantages: (1) it enables the model to handle tasks that are rarely encountered in the public corpus, and (2) it reduces the reliance on lengthy and complex prompts during inference, leading to faster processing times and lower costs [61].
Certain specialized tasks may require structured inputs and outputs. For instance, information extraction tasks or communication with physical devices require adherence to a specific input protocol and structured outputs, such as JSON objects, to ensure compatibility with machine parsing and processing [21]. Proprietary models often support features like JSON mode or function call capabilities [62], enabling such tasks to be handled efficiently through prompt optimization alone. Alternatively, open-source models can be fine-tuned on structured input-output pairs to achieve similar capabilities and can be further enhanced by employing constraint decoding frameworks [63]. Additionally, for tasks requiring up-to-date knowledge, such as summarizing findings from the latest clinical trials, RAG is essential. It dynamically incorporates external knowledge into the LLM inference process, ensuring that outputs remain current and relevant. For instance, GeneGPT [64] generates search queries to biomedical databases from the National Center for Biotechnology Information [65] to enhance the factual accuracy of biomedical question answering, achieving an accuracy of 0.83 compared to ChatGPT’s 0.12.
Evaluation is critical when adopting LLMs for medical applications, with the choice of evaluation metrics playing a central role. In many cases, general-domain metrics may not suffice. For example, in clinical note summarization tasks, general natural language processing metrics such as ROUGE [66] or BERTScore [67] may fail to adequately capture the quality of summaries in the medical domain. Instead, metrics tailored to the professional standards of medicine, such as completeness, correctness, and conciseness, should be prioritized [68]. Designing these domain-specific metrics is often a collaborative effort between AI researchers and medical practitioners, and it is advisable to consult relevant medical guidelines or regulatory policies to ensure the chosen metrics are appropriate and comprehensive. Furthermore, real-world user studies serve as the gold standard for evaluation. Such studies typically involve two arms, one leveraging AI-assisted workflows and the other relying on manual efforts, to assess the practical value and effectiveness of medical AI systems [6].